OpenStudy (anonymous):

What is the solution to the equation 64^(4-x)=4^(2x)? They want me to use complete sentences explaining how I did this so could you give me the steps on how to solve it? It has to be in approximate decimal form.

OpenStudy (shadowlegendx.):

hi

OpenStudy (anonymous):

hey. you wanna solve? lol. because I can't.

OpenStudy (shadowlegendx.):

What grade are you?

OpenStudy (anonymous):

11th

OpenStudy (shadowlegendx.):

I thought you were 19...

OpenStudy (anonymous):

it's a real long story, man. I had to drop out because of family and now im taking it online.

OpenStudy (shadowlegendx.):

oh....sorry for bringing it up >.>

OpenStudy (anonymous):

nah it's okay. you aren't gonna solve, are you?

OpenStudy (shadowlegendx.):

Whats the ^ mean?

OpenStudy (shadowlegendx.):

Powers?

OpenStudy (anonymous):

yeah, it's an exponent.

OpenStudy (anonymous):

I know more than you :3 that's sad... lol

OpenStudy (shadowlegendx.):

I technically already knew what it was...I was just making sure ;)

OpenStudy (anonymous):

oh okay, then by all means, proceed

OpenStudy (shadowlegendx.):

Idk...usually being called a turd doesnt inspire one to do work for ya ;)

OpenStudy (anonymous):

you don't even know how to do it, do you? lmao

OpenStudy (shadowlegendx.):

I probably do...somewhere in my head...but you should probably ask someone who hasnt been up for over a day ;)

OpenStudy (anonymous):

well I asked and you were the only who replied soo... lol. I took the ACT today so I have no more brain left for more math D:

OpenStudy (shadowlegendx.):

Get some rest then :)

OpenStudy (anonymous):

Psshh I wish. I just have this one assignment left and im trying to get it done

OpenStudy (shadowlegendx.):

Is it a " last minute " one ;P

OpenStudy (anonymous):

:/ yes. -____________-

OpenStudy (shadowlegendx.):

Welcome to the Club

OpenStudy (ranga):

64 = 4 x 4 x 4 = 4^3 Replace 64 with 4^3. Simplify. Since the bases are the same on both sides of the equation, the exponents must be the same. Equate the exponents and solve for x.

OpenStudy (shadowlegendx.):

Wow...nice was of putting it ranga

OpenStudy (anonymous):

Thank you, ranga, for helping me on this :) I appreciate it!

OpenStudy (shadowlegendx.):

You greens know your stuff :)

OpenStudy (ranga):

Thank you @ShadowLegendX. You are wleocme @megangray32.

OpenStudy (ranga):

\[\Large 64^{4-x} = 4^{2x}\]\[\Large (4^{3})^{4-x} = 4^{2x}\]\[\Large 4^{12 - 3x} = 4^{2x}\]Bases are the same. Therefore, exponents must be the same. \[\Large 12 - 3x = 2x\]Solve for x.

OpenStudy (anonymous):

12/5, right?

OpenStudy (ranga):

Yes, but they want the answer in decimal.

OpenStudy (anonymous):

oh yeah, it'd be 2.4 then?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

okay thank you again

OpenStudy (ranga):

Glad to be able to help.

OpenStudy (adrynicoleb):

*walks through question* Don't mind me, just stalking lalala~